OpenStudy (anonymous):

A ball is dropped from a height of 10m. It bounces to a height of 6m, then 3.6m, and so on following a geometric sequence. Find the total distance travelled by the ball. Hint: Use of sum to infinity and sum to certain term of a geometric sequence.

OpenStudy (anonymous):

who are you.

OpenStudy (anonymous):

first term = 10 common ratio = 0.6 now plug in the values in the formula

OpenStudy (anonymous):

sus*

OpenStudy (anonymous):

tell me the answer

OpenStudy (anonymous):

i know the first term is 10 and common ratio is 0.6 lol...

OpenStudy (anonymous):

my answer does not match with the answers at the back of the textbook

OpenStudy (anonymous):

what is the formula that you are using???

OpenStudy (anonymous):

Sum to infinity = a/(1-r)

OpenStudy (anonymous):

a being first term, r being common ratio

OpenStudy (anonymous):

Sum of geo series = a(1-r^n)/(1-r)

OpenStudy (anonymous):

a=10 (or not) r=0.6

OpenStudy (anonymous):

i got 25m, but the answers say 40m, proof pl0x

OpenStudy (anonymous):

thanks for viewing guys - need answers, any will do

OpenStudy (anonymous):

as long as it's correct

OpenStudy (anonymous):

Every term after the 10 is twice (down and up)...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

For 25 you counted once or twice?

OpenStudy (anonymous):

once, it's wrong

OpenStudy (anonymous):

so i guess it's something like 10+ ...sum from term 2 to infinity?

OpenStudy (anonymous):

i'll try it

OpenStudy (anonymous):

10 + twice * sum I think...

OpenStudy (anonymous):

a(1-r^n)/(1-r) = sum of geometric sequence to n a/(1-r) = sum to infinity 2x sum to infinity - first term!!!!!!! woot got it 25x2 - 10

OpenStudy (anonymous):

thanks :]

OpenStudy (anonymous):

i'll give you a medal for attempting

OpenStudy (anonymous):

tricky bastard questions... lol

OpenStudy (anonymous):

U r welcome.

OpenStudy (anonymous):

I will give u a medal for doing it, lol

OpenStudy (callisto):

Hmm... I can see there are 2..sum of inf. 1. first term =10, r=0.6 Sum = 10/(1-0.6) = 25 2. first term = 6, r=0.6 Sum = 6/(1-0.6) = 15 So.. Total sum = 25+15=40

OpenStudy (anonymous):

wow, i didn't see that

OpenStudy (callisto):

The first one is the one it drops... The 2nd one is the one it goes up. Total distance = total drops + total rise

OpenStudy (anonymous):

e^i pi ftw

OpenStudy (anonymous):

ahah!

OpenStudy (anonymous):

damn you're good

OpenStudy (callisto):

Do you get it now?

OpenStudy (anonymous):

yes, tyvm

OpenStudy (callisto):

You're welcome :)

OpenStudy (anonymous):

can you explain e^ipi plox?

OpenStudy (callisto):